A thin uniform, circular ring is rolling down an inclined plane of inclination 30° without slipping. Its linear acceleration along the inclined plane will be
(a) g/2
(b) g/3
(c) g/4
(d) 2g/3

Grade:12

1 Answers

Rs
6 years ago
a= gsin0 / (1+ mr^2)
 = g*(1/2)/(1+1) { i= mr^2} for ring
 = 9.8/2/2
=2.45 ans..............................................................................................................